OpenStudy (anonymous):

what should be added to both sides of this equation to solve for the variabel. 76= (Negative)-43+x

OpenStudy (ash2326):

we have \[\large 76=-43+x\] Add 43 to both the sides, tell me what do you get? @Gerber96

OpenStudy (anonymous):

119 and 86?

OpenStudy (ash2326):

If you add 43 , you'll get \[76+43=-43+43+x\] Notice there is a - before a 43 on the right side. So the two 43 will get cancelled \[76+43=\cancel{-43}\cancel{+43}+x\] What's x?

OpenStudy (anonymous):

is it 119 ?

OpenStudy (ash2326):

Yeah, do you understand?

OpenStudy (anonymous):

yes thank you